# A minimalistic and general weighted averaging method for inconsistent data

### v2

@@ Abstract
The weighted average of inconsistent data is a common and tedious problem that many scientists have encountered. The standard weighted average is not recommended for these cases, and different alternative methods are proposed in the literature. Here, we discuss a method first proposed by Sivia in 1996 that is based on Bayesian statistics and keeps the number of assumptions to a minimum. We propose this approach as a new standard for calculating weighted averages. The uncertainty associated with each input value is considered to be just a lower bound of the true unknown uncertainty. The resulting likelihood function is no longer Gaussian, but has smoothly decreasing wings, which allows for a better treatment of scattered data and outliers. The proposed method is tested on a series of data sets: simulations, CODATA recommended value of the Newtonian gravitational constant, and some particle properties from the Particle Data Group, including the proton charge radius. A freely available Python library is also provided for a simple implementation of the proposed averaging method.
@@ I Introduction
However, the inverse-invariance method, referred to in the following pages simply as standard, has a drawback. As we can see from Eq. (2), the final uncertainty depends only on the data uncertainties $\sigma_{i}$, but not on the data spread, which could be larger than the values of $\sigma_{i}$ (see, e.g., [1] for a more detailed discussion). This is, however, a common scenario in science, possibly caused by an uncontrolled systematic effect in the measurement procedure or by different biases in measurements conducted in different laboratories and/or with different methods. Common questions that arise are how to take into account such information on the data dispersion in the calculation of a weighted average and how to treat outliers.
To answer such questions, several approaches have been proposed in the literature. In the case of inconsistent data sets, a very common and basic method is to use the standard weighted average while artificially increasing its associated uncertainty. But how should one choose objectively the uncertainty expansion factor? The most common method has been proposed by Birge [2] almost one hundred years ago. It is based on the $\chi^{2}$ value obtained by the difference between the standard weighted average and the single input values. The uncertainty expansion factor $R_{\mathrm{Birge}}$, the Birge ratio, is applied to the single uncertainties $\tilde{\sigma}_{i}=R_{\mathrm{Birge}}\sigma_{i}$, with
@@
where $n$ is the number of $x_{i}$ data points. In this way, the final value of the reduced $\chi^{2}$ is adjusted to be close to unity, as expected for consistent data sets. The use of the Birge ratio is indeed one of the pillars of the statistical treatment employed by the Task Group on Fundamental Constants of the Committee on Data of the International Science Council (here simply abbreviated by CODATA) and the Particle Data Group (PDG). A modified version of the Birge ratio has been proposed in past works based on Bayesian statistics. The scaling factor $R$ between the estimated uncertainty $\sigma_{i}$ and the real uncertainty $\sigma_{i}^{\prime}$ is considered unknown but common to all data points [3, 4, 5, 1]. Assuming a non-informative Jeffreys’ prior probability $p(R)\propto 1/R$ and marginalising over the possible values of $R$, a final probability distribution is obtained with an average value (the mode of the final probability distribution) equal to the standard weighted average. The uncertainty corresponds to an expansion of the standard weighted average uncertainty by a factor equal to $R_{\mathrm{Bayes}}=\sqrt{(n-1)/(n-3)}R_{\mathrm{Birge}}$. Variations of this approach are discussed in Refs. [6, 7].
@@
In principle, because of the common scaling factor for each measurement result, the Birge ratio and its modified versions discussed above are not well suited for inter-laboratory averages, where very different systematic effects can occur. To compensate partially for such an issue, past works [3, 7] proposed to assign a random bias $\beta_{i}$ to each measurement, with a common mean value and standard deviation $\sigma_{\mathrm{bias}}$ for the entire ensemble of measurements. Here, a double marginalisation over the $\beta_{i}$ values and their shared uncertainty $\sigma_{\mathrm{bias}}$ is required, for which a prior probability has to be chosen, typically a non-informative Jeffreys’ prior. An evolution of such an approach has been proposed [8]. It consists of organising the input values into clusters, each with a different $\sigma_{\mathrm{bias}}$ value, and subsequently performing a Bayesian model average, in which each model corresponds to a different clustering choice.
An alternative approach, which like the Birge ratio avoids formulating any hypothesis on the nature of the missing systematic uncertainty, consists of estimating the uncertainty directly from the data dispersion with no particular assumption about the associated probability distribution [3, 9]. Datum-by-datum, the associated uncertainty $\tilde{\sigma}_{i}$ is obtained by a quadratic sum $\tilde{\sigma}_{i}=\sqrt{\sigma_{i}^{2}+d_{i}^{2}}$ of the known uncertainty $\sigma_{i}$ and the estimated missing uncertainty obtained from the difference $d_{i}=x_{i}-\hat{\mu}$ between the input value $x_{i}$ and the standard weighted average itself. Since $\hat{\mu}$ is thus present in both the left and right expressions of Eq. (1), the final average is obtained recursively. Like the original formulation of the Birge ratio, this method lacks statistical foundations but has the benefit of being very simply formulated.
@@
All the previously described methods share the implicit assumption that the uncertainty $\sigma_{i}$ is a lower bound of the real uncertainty $\sigma_{i}^{\prime}$. This simple and clear statement has been translated into formulas by Sivia and Skilling in 2004 [10], avoiding common scaling factors (like $R_{\mathrm{Birge}}$) or random bias dispersion (like $\beta_{\mathrm{bias}}$), but considering for each point a modification of the Gaussian distribution by the marginalisation over $\sigma_{i}^{\prime}$. For this approach, a prior probability $p(\sigma_{i}^{\prime})$ for $\sigma_{i}^{\prime}$ has to be chosen. The natural choice would be the non-informative Jeffreys’ prior $p(\sigma_{i}^{\prime})\propto 1/\sigma_{i}^{\prime}$. If not constraint by an upper bound, this choice causes divergence because of the non-integrability of the resulting final probability distribution associated to each datum. To avoid this problem, a modified prior $p(\sigma_{i}^{\prime})\propto 1/(\sigma_{i}^{\prime})^{2}$ has been proposed and discussed in Refs. [10, 11, 6]. Alternatively, the inverse-gamma distribution, which is integrable but introduces additional parameters, is proposed as a prior distribution [1, 12]. Other more complex approaches with no lower bounds for $\sigma_{i}^{\prime}$ can be found in Refs. [10, 1].
@@
Another approach involves the assumption of a likelihood that is not Gaussian. For example, Dose [13] proposes using Laplace and hyperbolic cosine likelihoods in addition to the Gaussian likelihood and averaging the three distributions by applying Bayesian model selection.
@@
Here, we consider an approach with very few assumptions, similar to the work of Sivia and Skilling [10], but adopt the Jeffreys’ non-informative prior, like in Sivia’s earlier work [14]. Our intent is not to propose an automatic approach to replace the work of data selection, but to offer an easily implementable method as a common reference for averaging, while avoiding the issues associated with the standard weighted average. For these reasons, like the standard weighted average, we keep the basic assumptions simple and very general, with the risk of being too conservative in the estimation of the final uncertainty in certain cases. To facilitate the application of the proposed method in daily data analysis, a Python code is provided.
@@ II.1 General considerations
The standard weighted average of independent measurements $x_{i}$ with uncertainties $\sigma_{i}$ is obtained by maximising the total probability of the mean value $\mu$, which results from the product of the single probabilities for each (assumed independent) measurement, with
$p(\mu|\{x_{i},\sigma_{i}\})=\prod_{i}p(x_{i}|\mu,\sigma_{i})\ p(\mu).$ (4)
When a Gaussian distribution is considered for each $x_{i}$ together with a flat prior probability for $\mu$ (a Jeffreys’ prior), the most probable value $\hat{\mu}$ is given by the standard weighted average, i.e., Eq. (1). The associated uncertainty $\sigma_{\hat{\mu}}$ given in Eq. (2) is simply derived by the uncertainty propagation in $\hat{\mu}=f(x_{1},x_{2},\ldots)$ of the uncertainty of the single $x_{i}$ values (see e.g. [15, 16]).
@@
In line with standard methods, we consider the average as the value $\hat{\mu}$ that maximises Eq. (4) (assuming the Jeffreys’ prior $p(\mu)=\text{const.}$.), with its uncertainty given by Eq. (5). The single probabilities $p(x_{i}|\mu,\sigma_{i})$ are no longer Gaussian, but are instead obtained by assuming certain hypotheses on the priors and performing marginalisations.
As wrote above, our final goal is to provide the scientific community with a new accessible tool to obtain a robust weighted average that can be easily understood. Two prerogatives are thus essential: to propose something very general and to consider a minimal number of assumptions. The generality is particularly important to treat very common but different scenarios of inconsistent data averaging: i) from measurements obtained with the same apparatus (with a common uncontrolled systematic effect) or ii) from different types of measurements in different laboratories (with possible uncorrelated biases). For these purposes, we adopt a pessimistic framework where the uncertainty $\sigma_{i}$ is regarded as a lower bound of the real uncertainty $\sigma^{\prime}_{i}$, without any additional assumptions on the possible biases and relations influencing the available data set (similarly to Refs. [10, 17]). Any systematic error is considered to be included in the unknown uncertainty $\sigma_{i}^{\prime}$.
@@
If only pairs $(x_{i},\sigma_{i})$ of measured values and associated uncertainties are available, following the maximum entropy principle, a Gaussian distribution $p(x_{i}|\mu,\sigma_{i})$ can be assumed for each datum. A choice for the prior probability distribution $p(\sigma^{\prime}_{i}|\sigma_{i})$ for $\sigma^{\prime}_{i}$ has to be made. The natural choice is a Jeffreys’ prior, which is a non-informative prior that is invariant under reparametrisation, to avoid introducing other possible biases, with
$\begin{cases}p(\sigma^{\prime}_{i}|\sigma_{i})=\cfrac{1}{\log(\sigma^{\mathrm{max}}_{i}/\sigma_{i})}\cfrac{1}{\sigma_{i}^{\prime}}&\text{\quad for }\ \sigma^{\prime}_{i}\leq\sigma_{i}\leq\sigma^{\mathrm{max}}_{i},\\ 0&\text{\quad otherwise.}\\ \end{cases}$ (7)
The problem of such a prior is the introduction of an additional parameter $\sigma^{\mathrm{max}}_{i}$ for each data point. Indeed, if one tries to eliminate such additional parameters by considering the limit $\sigma^{\mathrm{max}}_{i}\to\infty$, $p(\sigma^{\prime}_{i}|\sigma_{i})$ is no longer a proper probability distribution, i.e. normalised to unity, because $\int_{\sigma_{i}}^{\infty}1/\sigma^{\prime}_{i}\mathrm{d}\sigma^{\prime}_{i}=\infty$. Two alternative solutions to this issue are presented in the next paragraphs.
@@ II.2 Sivia and Skilling’s conservative weighted average
To avoid the introduction of additional parameters $\sigma^{\mathrm{max}}_{i}$ required for the normalisation of Jeffreys’ prior, a conservative formulation has been proposed by Sivia and Skilling for general regression problems [10]. It consists of the modified version of the Jeffreys’ prior
@@
Keeping the assumption of a Gaussian distribution for $p(x_{i}|\mu,\sigma^{\prime}_{i})$ and combining Eqs. (6) and  (8), we obtain
$p(x_{i}|\mu,\sigma_{i})=\frac{\sigma_{i}}{\sqrt{2\pi}}\left[\frac{1-e^{\frac{(x_{i}-\mu)^{2}}{2\sigma^{2}_{i}}}}{(x_{i}-\mu)^{2}}\right].$ (9)
@@ II.3 Limit solution with Jeffreys’ prior
A criticism that could be directed at Eq. (9) is that the choice of the prior probability distribution of $\sigma_{i}^{\prime}$ in Eq. (8) does not respect the non-informative criterion of Jeffreys’ prior. A possible solution to keep Jeffreys’ prior without introducing any new parameters is given by Sivia in Ref. [14]. It considers the limit case $\sigma^{\mathrm{max}}_{i}\to\infty$. The divergence of the resulting probability distribution is circumvented by considering only its maximum and the second-order log-derivative for the estimation of the limit value of the weighted average and the associated uncertainty, respectively.
When the Jeffreys’ prior from Eq. (7) is adopted, Eq. (6) becomes
@@
Compared to Eq. (9), we can note that the distribution tails decrease even more smoothly than in the conservative approach, with a dependency of $1/x_{i}$ instead of $1/x_{i}^{2}$, which can tolerate the presence of outliers even better. These more pronounced tails are well visible in Fig. 1. The logarithmic form of the total probability $p(\mu|\{x_{i},\sigma_{i},\sigma^{\mathrm{max}}_{i}\})$ is thus given by
@@
where the constant $C^{\infty}=\lim_{\sigma^{\mathrm{max}}_{i}\to\infty}C=\infty$ is indeed divergent, but for more than one data point the distribution is integrable, and the variance is finite for more than two data points. In particular, the position of the maximum and the value of the second derivative and thus the weighted average and its uncertainty are well defined. As in the case of the conservative weighted average, no analytical solution is available for $\hat{\mu}$ and $\sigma_{\hat{\mu}}$, so the solution must be found numerically.
We consider the formulas presented in this section as a new reference method for the weighted average. In fact, the conservative approach of Sec. II.2 still involves some arbitrariness in the choice of the prior for the average. This is not the case for the above formulas, where the universal choice of the Jeffreys’ prior is adopted. In the following section, we will refer to this approach as the Jeffreys’ weighted average.
@@ III Some applications
In this section, we present a series of applications for common data analysis cases. In the first subsection, we will study simulated data with known theoretical values of mean and standard deviation. In the second subsection, an analysis of the different values of the Newtonian gravitational constant from past CODATA compilations is proposed. The third subsection is dedicated to fundamental particle properties, including the controversial data set of the proton charge radius. Additional calculations of the Jeffreys’ weighted average for inconsistent data measured using the same experimental apparatus can be found in Ref. [18].

### v3

# A simple tool for weighted averaging of inconsistent data sets

@@ Abstract
The weighted average of inconsistent data is a common and tedious problem that many scientists have encountered. The standard weighted average is not recommended for these cases, and various alternative methods have been proposed. These approaches vary in suitability depending on the nature of the data, which can make selecting the appropriate method difficult without expertise in metrology or statistics. For the analysis of simple data sets presenting inconsistencies, we discuss the method proposed by Sivia in 1996 based on Bayesian statistics. This choice has the intention of maintaining generality while minimising the number of assumptions. In this approach, the uncertainty associated with each input value is considered to be just a lower bound of the true unknown uncertainty. The resulting likelihood function is no longer Gaussian but has smoothly decreasing wings, which allows for a better treatment of scattered data and outliers. To demonstrate the robustness and the generality of the method, we apply it to a series of critical data sets: simulations, CODATA recommended values of the Newtonian gravitational constant, and some particle properties from the Particle Data Group, including the proton charge radius. A freely available Python library is also provided for a simple implementation of the proposed averaging method.
@@ I Introduction
However, the inverse-invariance method, referred to in the following pages simply as standard, has a drawback. As we can see from Eq. (2), the final uncertainty depends only on the data uncertainties $\sigma_{i}$, but not on the data spread, which could be larger than the values of $\sigma_{i}$ (see, e.g., vonderLinden for a more detailed discussion). This is, however, a common scenario in science, possibly caused by an uncontrolled systematic effect in the measurement procedure or by different biases in measurements conducted in different laboratories and/or with different methods. Common questions that arise are how to take into account such information on the data dispersion in the calculation of a weighted average and how to treat outliers.
To answer such questions, several approaches have been proposed in the literature. A very common and basic method is to use the standard weighted average while artificially increasing its associated uncertainty. But how should one choose objectively the uncertainty expansion factor? The most common method was proposed by Birge Birge1932 almost one hundred years ago. It is based on the $\chi^{2}$ value obtained by the difference between the standard weighted average and the single input values. The uncertainty expansion factor $R_{\mathrm{Birge}}$, the Birge ratio, is applied to the single uncertainties $\tilde{\sigma}_{i}=R_{\mathrm{Birge}}\sigma_{i}$, with
@@
where $n$ is the number of $x_{i}$ data points. In this way, the final value of the reduced $\chi^{2}$ is adjusted to be close to unity, as expected for consistent data sets. A modified version of the Birge ratio has been proposed in past works based on Bayesian statistics. The scaling factor $R$ between the estimated uncertainty $\sigma_{i}$ and the real uncertainty $\sigma_{i}^{\prime}$ is considered unknown but common to all data points Lira2007; Toman2012; Bodnar2014; vonderLinden. Assuming this, along with a reasonable prior on the value of $R$, the modified expansion factor $R_{\mathrm{Bayes}}=\sqrt{(n-1)/(n-3)}R_{\mathrm{Birge}}$ is obtained. Variations of this approach are discussed in Refs. Mana2012; Bodnar2016.
@@
Because of the common scaling factor for each measurement result, the Birge ratio and its modified versions discussed above are however not well suited for interlaboratory averages, where very different systematic effects can occur. To compensate partially for such an issue, past works Lira2007; Bodnar2016 proposed to assign a random bias $\beta_{i}$ to each measurement, with a common mean value and standard deviation $\sigma_{\mathrm{bias}}$ for the entire ensemble of measurements. Here, a double marginalisation over the $\beta_{i}$ values and their shared uncertainty $\sigma_{\mathrm{bias}}$ is required. An evolution of such an approach has been proposed Rukhin2019.
@@
Many other approaches exist (see, e. g., vonderLinden; Dose2003; Possolo2023 for a general overview). Like the methods discussed above, they implicitly assume that the uncertainty $\sigma_{i}$ is a lower bound of the real uncertainty $\sigma_{i}^{\prime}$. This simple and clear statement has been translated into formulas by Sivia and Skilling Sivia1996b; Sivia, avoiding common scaling factors (like $R_{\mathrm{Birge}}$) or random bias dispersion (like $\beta_{\mathrm{bias}}$), but considering for each point a modification of the Gaussian distribution by the marginalisation over $\sigma_{i}^{\prime}$. For this approach, a prior probability $p(\sigma_{i}^{\prime})$ for $\sigma_{i}^{\prime}$ has to be chosen. The natural choice would be the non-informative Jeffreys’ prior $p(\sigma_{i}^{\prime})\propto 1/\sigma_{i}^{\prime}$. This prior is indeed invariant over a change of coordinates for the parameter of interest with a minimization of possible biases. Its implementation is especially interesting for scale parameters such as $\sigma_{i}$. (See e.g. Ref. vonderLinden for a detailed discussion.) If not constrained by an upper bound, this choice causes divergence because of the non-integrability of the resulting final probability distribution associated with each datum. To avoid this problem, a modified prior $p(\sigma_{i}^{\prime})\propto 1/(\sigma_{i}^{\prime})^{2}$ has been proposed and discussed in Refs. Sivia; Shirono2010; Mana2012. Alternatively, the inverse-gamma distribution, which is integrable but introduces additional parameters, is proposed as a prior distribution vonderLinden; Dose2003. Other more complex approaches with no lower bounds for $\sigma_{i}^{\prime}$ can be found in Refs. Sivia; vonderLinden.
@@
Note that alternative approaches based on non-Gaussian distributions have also been discussed in the literature. In particular, the Student’s t-distribution is well known for its robustness to outliers Dawid1973; OHagan1979. Dose Dose2007 proposes using Laplace and hyperbolic cosine likelihoods in addition to the Gaussian likelihood and averaging the three distributions by applying Bayesian model selection.
@@
Here, we discuss in detail the method by Sivia and Skilling Sivia1996b; Sivia. As will be seen, this approach is easily implementable while avoiding the issues associated with the standard weighted average. Like the standard weighted average, basic assumptions are kept simple and very general. To facilitate the application of the proposed method in daily data analysis, a Python code is provided.
@@ II.1 General considerations
The standard weighted average of independent measurements $x_{i}$ with uncertainties $\sigma_{i}$ is obtained by maximising the total probability $p(\mu|\{x_{i},\sigma_{i}\})$ of the mean value $\mu$ for a given set of measured values $\{x_{i},\sigma_{i}\}$. Because of the independence of the measurements, it results from the product of the single probabilities for each measurement, with
$p(\mu|\{x_{i},\sigma_{i}\})\propto\prod_{i}p(x_{i}|\mu,\sigma_{i})\ p(\mu).$ (4)
With the above formula, we implicitly assume that the probability distribution of each datum $x_{i}$ can be described by a distribution with only two parameters, its mean and standard deviation, and that the best estimation of the standard deviation is given by the provided values $\sigma_{i}$. In the following sections, we will question the latter assumption.
When a Gaussian distribution is considered for each $x_{i}$ together with a flat prior probability for $\mu$ (a Jeffreys’ prior), the most probable value $\hat{\mu}$ is given by the standard weighted average, i.e., Eq. (1). The associated uncertainty $\sigma_{\hat{\mu}}$ given in Eq. (2) is simply derived by the uncertainty propagation in $\hat{\mu}=f(x_{1},x_{2},\ldots)$ of the uncertainty of the single $x_{i}$ values (see e.g. Bevington; Taylor).
@@
In line with standard methods, we consider the average as the value $\hat{\mu}$ that maximises Eq. (4) (assuming the Jeffreys’ prior $p(\mu)=\text{const.}$), with its uncertainty given by Eq. (5). The single probabilities $p(x_{i}|\mu,\sigma_{i})$ are no longer Gaussian, but are instead obtained by assuming certain hypotheses on the priors and performing marginalisations.
As stated above, our ultimate goal is to provide a tool for obtaining a robust weighted average applicable to a broad range of cases, which can be easily used as an alternative to the standard inverse-variance weighted average. Two prerogatives are thus essential: to propose something very general and to consider a minimal number of assumptions. For these purposes, we adopt the pessimistic framework where the uncertainty $\sigma_{i}$ is regarded as a lower bound of the real uncertainty $\sigma^{\prime}_{i}$, without any additional assumptions on the possible biases and relations influencing the available data set. Any systematic error is considered to be included in the unknown uncertainty $\sigma_{i}^{\prime}$.
@@
If only pairs $(x_{i},\sigma_{i})$ of measured values and associated uncertainties are available, a Gaussian distribution $p(x_{i}|\mu,\sigma_{i})$ can reasonably be assumed for each datum. A choice for the prior probability distribution $p(\sigma^{\prime}_{i}|\sigma_{i})$ for $\sigma^{\prime}_{i}$ has to be made. The natural choice is a Jeffreys’ prior, which is a non-informative prior that is invariant under reparametrisation to avoid introducing other possible biases, with
[…]
The problem of such a prior is the introduction of an additional parameter $\sigma^{\mathrm{max}}_{i}$ for each data point.
@@ II.2 Sivia and Skilling’s conservative weighted average
To avoid the introduction of additional parameters $\sigma^{\mathrm{max}}_{i}$ and keep a proper probability distribution for $p(\sigma^{\prime}_{i}|\sigma_{i})$, an alternative prior has been proposed by Sivia and Skilling Sivia:
@@
We will refer in the following section to this approach as the conservative approach, consistent with the authors’ own terminology. Keeping the assumption of a Gaussian distribution for $p(x_{i}|\mu,\sigma^{\prime}_{i})$ and combining Eqs. (6) and  (8), we obtain
$p(x_{i}|\mu,\sigma_{i})=\frac{\sigma_{i}}{\sqrt{2\pi}}\left[\frac{1-e^{-\frac{(x_{i}-\mu)^{2}}{2\sigma^{2}_{i}}}}{(x_{i}-\mu)^{2}}\right].$ (9)
@@ II.3 Limit solution with Jeffreys’ prior
A criticism that could be directed at Eq. (9) is that the choice of the prior probability distribution of $\sigma_{i}^{\prime}$ in Eq. (8) does not respect the non-informative criterion of Jeffreys’ prior. A possible solution to keep Jeffreys’ prior without introducing any new parameters is given by Sivia in Ref. Sivia1996b. It considers the limit case $\sigma^{\mathrm{max}}_{i}\to\infty$. While this approach can be considered even more conservative than the previous one, we will refer to it in the following section as Jeffreys’ prior approach.
When the Jeffreys’ prior from Eq. (7) is adopted and $\sigma^{\mathrm{max}}_{i}$ have a finite value, Eq. (6) becomes
@@
Compared to Eq. (9), we can note that the distribution tails decrease even more smoothly than in the conservative approach, with a dependency of $1/x_{i}$ instead of $1/x_{i}^{2}$, which can tolerate the presence of outliers even better. These more pronounced tails are well visible in Fig. 1. For each individual datum, $p(x_{i}|\mu,\sigma_{i})$ is an improper probability distribution because it is non-integrable. However, when at least two data values are considered, the probability $p(\mu|\{x_{i},\sigma_{i}\})$ (Eq. (4)) is well defined, and the weighted average can be determined without ambiguities.
The logarithmic form of the total probability $p(\mu|\{x_{i},\sigma_{i},\sigma^{\mathrm{max}}_{i}\})$ is given by
@@
where the constant $C^{\infty}=\lim_{\sigma^{\mathrm{max}}_{i}\to\infty}C=\infty$ is indeed divergent, but for more than one data point, the distribution is integrable, and the variance is finite for more than two data points. In particular, the position of the maximum and the value of the second derivative and thus the weighted average and its uncertainty are well defined. As in the case of the conservative weighted average, no analytical solution is available for $\hat{\mu}$ and $\sigma_{\hat{\mu}}$, so the solution must be found numerically.
In the following section, we will refer to this approach as the Jeffreys’ weighted average.
@@ III Some applications
In this section, we present a series of applications for common data analysis cases to test the reliability of the conservative and Jeffreys’ weighted average. In the first subsection, we will study simulated data with known theoretical values of mean and standard deviation. In the second subsection, an analysis of the different values of the Newtonian gravitational constant from past CODATA compilations is proposed. The third subsection is dedicated to fundamental particle properties, including the controversial data set of the proton charge radius. Additional implementations of the Jeffreys’ weighted average for inconsistent data measured using the same experimental apparatus can be found in Ref. Duval2024.